Having trouble. Can you help? Evaluate (f o g)(3), given that f(x)=x^2-5x+1 and g(x)= sqrt(x+1). Thank you.
Answer by stanbon(48535)

Evaluate (f o g)(3), given that f(x)=x^2-5x+1 and g(x)= sqrt(x+1)
-----------
f[g(x)] = f[sqrt(x+1)]
= [sqrt(x+1)]^2-5[sqrt(x+1)]+1
= x+1-5sqrt(x+1) + 1
= x+2-5sqrt(x+1)
===============
Therefore (fog)(3) = 3+2-5sqrt(4)= 3+2-5*2 = -5
=================
Cheers,
Stan H.